OpenStudy (anonymous):

3x^2 + 4x + 5x^2 - 5y

OpenStudy (anonymous):

what is the question ?

OpenStudy (anonymous):

\[\LARGE ax^2+bx^2=(a+b)x^2\]

OpenStudy (anonymous):

\[\LARGE x^2 = \text{apples}\] \[\LARGE 3\;\; \text{apples}+5\;\;\text{apples}= \text{how many apples ?}\]

OpenStudy (anonymous):

simplify the equation

OpenStudy (linna):

LOL